The equilibrium constant Kc for the reaction H2(g) + Br2(g) ⇆ 2HBr(g) is 2.180 × 106 at 730°C. Starting with 1.20 moles of HBr in a 21.3−L reaction vessel, calculate the concentrations of H2, Br2, and HBr at equilibrium.
Answer #1

Solution е анви са ) H, 9) + Вися) K = 9, 10 Time = +30°C V= 21.3L Inital moles of HBu = 1.20 moles. моаси +: 1:20 &1. 3 o-o510 х 10 + 2 = o-o5 6 и х» з 14 x 5 (s) [н,= 3-774 х 10 см ) [BA.] : 3. a21 хом [НВА) , 0.656 - 9 - б•o 5. - 2x (3. 2 x 10 ° )please hit thumbs up this question. .if u like it .....and keep posting your doubts ... thank you so much